# SMILES has to go : Representation of Molecules via Algebraic Data Types

### v1

@@ 2.4 Molecules and programming languages
Inductive Logic Programming (ILP), has seen usage in representing basic molecules Srinivasan et al. [1996], but the focus was doing rule induction over graphical data structures, with molecules being an example, rather than presenting ILP as a representational format for cheminformaticians. It is perhaps for this reason that it has had little uptake within the Cheminformatics community.
Molecular representations using Inductive Logic Programming (ILP) have been explored in previous research, such as the work by Srinivasan et al. [Srinivasan et al., 1996]. However, the application of machine learning techniques in conjunction with ILP-based molecular representations has not been extensively discussed in this early work. This lack of emphasis on machine learning aspects may explain the absence of the paper from recent reviews of molecular representations, such as the comprehensive survey by Wigh et al. [Wigh et al., 2022].
In the approach presented by Srinivasan et al. [Srinivasan et al., 1996], molecules are represented as logical rules using an extension of the Prolog programming language. Prolog is a declarative language that enables the expression of logical relationships between variables using first-order logic. Programs written in Prolog consist of a complex network of facts and rules, which are constructed using logical primitives such as conjunction, disjunction, and universal quantification. However, the data types supported by Prolog are limited to lists, integers, floats, and strings, with minimal support for user-defined data types.
@@
The specific extension of Prolog used in [Srinivasan et al., 1996] is called Progol, which employs inductive logic programming techniques to derive rules that describe the given data. It is worth clarifying that Progol should not be confused with ProbLog, a probabilistic logic programming language that focuses on modeling the probabilities of examples being true based on the probabilities of the underlying relations.
@@
While the work by Srinivasan et al. [Srinivasan et al., 1996] demonstrates the feasibility of representing molecules using logical rules in an ILP framework, the integration of machine learning techniques with these representations was not explored.
This makes the proposal here the first to suggest to cheminformaticians to use programming languages more directly in representation. The notion that “molecular programming languages” could be possible, occurred in an article about the future of SELFIES [Krenn et al., 2022]. It is mentioned that molecular graphs could be executed, but the language is ambiguous and short. It is unclear whether they suggest using molecules to represent programming languages via the Church-Turing thesis, whether they want to compute with graph structures or are attempting to compute in a substrate other than silicon  [^3] https://molecularprogrammers.org/.
@@
Investigations of type safe functional programming for cheminformatics has occurred before. Recent work has suggested the use of OCaml in writing software for cheminformatics [Berenger et al., 2019], [Höck and Riedl, 2012]. However, the emphasis in both of these articles as well as the Haskell library Ouch [T, 2021], was not on extensible molecular representations but on functional programming. This is highlighted by the papers introduction of lengthy tutorials on functional programming. Furthermore they do not show how Bayesian Inference can be performed on drug structures over algebraic data types which is essential in motivating efficiency improvements in not requiring exploration of syntactically nonsensical regions of molecular space.
2.4.1 RDKit as representation
One might also ask whether Cheminformatics libraries support representation. Behemoth Cheminformatics programming libraries, such as RDKit, [Landrum et al., 2013] and lightweight molecular manipulation libraries that exist on top of these e.g. datamol [^4] https://datamol.io/ and https://www.rdkit.org/ (which builds on NumPY [Oliphant et al., 2006], Pandas, RDKit and Python 3+), Datamol explicitly [^5] https://doc.datamol.io/stable/tutorials/Preprocessing.html don’t advocate themselves as providing a prototypical representation. These libraries parse representations to computational representations and then apply functions to them, before outputting them to strings like SMILES.
The computational representation of RDKit in C++ is complex, taking up over 400 lines of highly complex code and the molecule class taking up 873 lines of code. There are no clear invariants, or guarantees about the management of side effects in snippets shown below. The ubiquitous prescence of null pointers and highly abstruse C++ keywords, scattered throughout the code, insinuates points of failure of the program. It is unclear, whether and when the program may crash and what the consequences might be. The following snapshot of lines $792$ to $821$ in ROMol.h [^6] https://www.rdkit.org/docs/cppapi/ROMol_8h_source.html which represents the molecule class, provides a brief glimpse into the extreme complexity of code:
Taken from https://www.rdkit.org/docs/cppapi/ROMol_8h_source.html
--Line 792
    #ifdef RDK_USE_BOOST_SERIALIZATION
[…]
   template <class Archive>
   void save(Archive &ar, const unsigned int version) const;
   template <class Archive>
   void load(Archive &ar, const unsigned int version);
   BOOST_SERIALIZATION_SPLIT_MEMBER()
 #endif
  private:
   MolGraph d_graph;
   ATOM_BOOKMARK_MAP d_atomBookmarks;
   BOND_BOOKMARK_MAP d_bondBookmarks;
   RingInfo *dp_ringInfo = nullptr;
   CONF_SPTR_LIST d_confs;
   std::vector<SubstanceGroup> d_sgroups;
   std::vector<StereoGroup> d_stereo_groups;
   std::unique_ptr<boost::dynamic_bitset<>> dp_delAtoms = nullptr;
   std::unique_ptr<boost::dynamic_bitset<>> dp_delBonds = nullptr;
   friend RDKIT_GRAPHMOL_EXPORT std::vector<SubstanceGroup> &getSubstanceGroups(
       ROMol &);
   friend RDKIT_GRAPHMOL_EXPORT const std::vector<SubstanceGroup>
       &getSubstanceGroups(const ROMol &);
   void clearSubstanceGroups() { d_sgroups.clear(); }
  protected:
   unsigned int numBonds{0};
-- Line 821
Figure 4: This is just a small fragment of the molecule class in RDKit. It demonstrates the sheer complexity and obfuscation of a concept that can be represented more simply in the ADT proposed here.
The proposed ADT-based representation of molecules in Haskell offers a more concise, expressive, and type-safe approach. By leveraging the power of algebraic data types and the strong static typing of Haskell, we can define molecular structures in a declarative and compositional manner. The resulting code is more readable, maintainable, and less prone to runtime errors. The use of pure functions and the absence of side effects in the ADT representation enable easier reasoning about the behavior of the program and facilitate formal verification techniques.
Moreover, the ADT-based representation provides a clear separation of concerns between the data structure and the operations performed on it. This modular design allows for greater flexibility and extensibility, as new operations can be defined independently of the underlying data representation. The use of higher-order functions and type classes in Haskell further enhances the reusability and composability of the code.
While cheminformatics libraries like RDKit and datamol have been widely adopted and have proven their utility in various applications, their complex OOP-based representations can hinder the development and maintenance of robust and reliable software. The proposed ADT-based approach offers a compelling alternative that prioritizes simplicity, type safety, and mathematical rigor.
2.5 Probabilistic Programming
Languages that are said to be probabilistic are defined as those that support both random sampling from distributions and the conditioning of values of a variable via observations [Gordon et al., 2014]. At the core of probabilistic programming are two essential operations: sampling and scoring. Sampling allows for the generation of random values from probability distributions, while scoring enables the assignment of likelihoods to observed data given a probabilistic model. These operations form the foundation for performing Bayesian inference, where the goal is to infer the posterior distribution of model parameters given observed data. Furthermore, if programs featuring the ADT are written in do notation, as the probability probability / measure monad is commutative [Staton, 2017], consecutive sampling statements can be done independently and in parallel.
In the proposed approach, the Molecule algebraic data type (ADT) serves as the basis for constructing a generative model of molecules. By representing molecules as structured data using the ADT, we can define probability distributions over various molecular properties, such as atom types, bond types, and spatial coordinates.
We can define probability distributions over other molecular properties, such as bond types and electron configurations, using appropriate probability distributions.
Given a dataset of observed molecules, we can perform Bayesian inference to estimate the posterior distribution of the model parameters. One popular inference algorithm is Trace Metropolis-Hastings (Trace-MH), which is a Markov Chain Monte Carlo (MCMC) method adapted for probabilistic programming.
Trace-MH operates by proposing modifications to the execution trace of the probabilistic program, which represents a particular instantiation of the model parameters and random variables. The proposed trace is then accepted or rejected based on the Metropolis-Hastings acceptance criterion, which compares the likelihood of the proposed trace to the current trace.
Furthermore, the use of probabilistic programming allows for the incorporation of prior knowledge and the quantification of uncertainty in molecular modeling. By specifying informative priors over molecular properties and parameters, domain expertise can be incorporated into the modeling process. Additionally, the posterior distributions obtained through Bayesian inference provide a measure of uncertainty associated with the inferred molecular structures and properties, enabling more robust and reliable predictions.
The proposed approach also facilitates the integration of probabilistic programming with other machine learning techniques, such as deep learning. By leveraging the automatic differentiation capabilities of modern deep learning frameworks, it becomes possible to train complex neural network models within the probabilistic programming framework. This enables the development of hybrid models that combine the expressiveness of deep learning with the principled handling of uncertainty provided by probabilistic programming.
The proposed approach of using algebraic data types (ADTs) to represent molecules offers a significant advantage in terms of efficiency and meaningful exploration of chemical space. By directly synthesizing molecules in accordance with the grammar defined by the ADTs, the machine learning process can operate directly on the program or grammar itself, without the need for distilling the representation into another format.
This direct operation on the grammatical representation has important implications for the efficiency and effectiveness of the machine learning process. One of the key benefits is that it eliminates the need to explore nonsensical or invalid regions of chemical space. The grammar inherently constrains the space of possible molecules to those that are semantically meaningful and consistent with the rules and structure defined by the ADTs.
In traditional approaches, where molecules are represented using string-based formats like SMILES or graph-based representations, the machine learning algorithms often generate or explore a large number of candidate molecules, many of which may be invalid or chemically infeasible. This exploration of nonsensical regions of chemical space is computationally wasteful and can hinder the efficiency and effectiveness of the learning process.
By operating directly on the grammatical representation, the machine learning algorithm can focus its exploration on the space of valid and meaningful molecules. The ADTs provide a structured and constrained representation that ensures the generated molecules adhere to the specified rules and constraints. This targeted exploration of the chemical space leads to more efficient and meaningful learning, as the algorithm can avoid wasting computational resources on exploring invalid or nonsensical molecules.
Furthermore, the direct operation on the grammar allows for the seamless integration of domain knowledge and chemical constraints into the learning process. The ADTs can be designed to incorporate specific rules, symmetries, and constraints that are relevant to the problem at hand. By encoding this knowledge directly into the representation, the machine learning algorithm can leverage this information to guide its search and generate molecules that are more likely to exhibit desired properties or meet specific criteria.
The probabilistic conditioning operations, such as sampling and scoring, can be directly applied to the grammatical representation, enabling the algorithm to assign probabilities and likelihoods to different molecular structures and properties. This probabilistic treatment allows for the quantification of uncertainty and the incorporation of prior knowledge into the learning process.
Moreover, the use of a grammatical representation facilitates the interpretability and explainability of the generated molecules. Since the molecules are synthesized directly from the grammar, they inherently possess a structured and meaningful representation. This enables chemists and domain experts to easily interpret and understand the generated molecules, as they align with the defined rules and constraints of the ADTs.
In contrast, when molecules are distilled into other representations, such as fixed-length vectors or arbitrary graph structures, the interpretability and chemical meaning of the generated molecules can be lost or obscured. The direct operation on the grammar preserves the semantic structure and allows for a more intuitive understanding of the generated molecules.
One of the key benefits of using probabilistic programming for retrosynthetic analysis is the ability to directly generate and manipulate molecular graphs or trees. In traditional approaches, molecules are often represented using string-based formats like SMILES or graph-based representations, which require additional preprocessing and conversion steps. However, with probabilistic programming languages like LazyPPL, molecules can be directly synthesized as algebraic data types (ADTs) that adhere to a well-defined grammar. This allows for the direct generation and manipulation of molecular structures without the need for intermediate representations.
Previous implementations of Gaussian Processes have been restricted by the grammar [Moss and Griffiths, 2020]. By using Gaussian processes in combination with probabilistic programming, we can model the probability distributions over molecular properties and reactivity. Gaussian processes provide a flexible and expressive framework for capturing the complex relationships between molecular features and their corresponding properties. They allow for the incorporation of prior knowledge and the quantification of uncertainty in the predictions.
In the context of retrosynthetic analysis, Gaussian process regression can be used to predict the likelihood of a particular synthetic route or the feasibility of a retrosynthetic step. By training the Gaussian process model on a dataset of known reactions and their outcomes, it can learn the underlying patterns and relationships between molecular structures and reactivity. This learned model can then be used to guide the exploration of the chemical space and prioritize promising retrosynthetic pathways.
The integration of probabilistic programming with retrosynthetic analysis enables a more efficient and targeted exploration of the chemical space. By directly growing molecular trees or graphs and subjecting them to probabilistic conditioning operations, we can focus on the most relevant and plausible regions of the chemical space. The scoring and sampling mechanisms provided by probabilistic programming languages allow for the assignment of probabilities to different retrosynthetic routes and the selection of the most promising candidates.
2.6 Introduction to the proposal
The proposed representation here, is embedded in a pure, type-safe, functional programming language called Haskell. This purity makes reasoning about, and debugging code easier. Types, together with type inference allows one to have correctness guarantees at compile-time, eliminating runtime errors. None of the concepts presented here are conceptually specific to Haskell and are found in other functional programming languages.
The proposed representation is also motivated by established and the conceptually well motivated interest of Bayesian machine learning for drug design [Moss and Griffiths, 2020], [Semenova et al., 2020] & [Mervin et al., 2021]. Integration of probabilistic programming with Haskell e.g. [Paquet and Staton, 2021], [Ścibior et al., 2015], [Nguyen et al., 2022] has been done before laying the groundwork for usage with chemical data structures.
Algebraic Data Types (ADTs) provide a natural and intuitive way to represent molecular structures because they implicitly define a grammar for the data. ADTs are composed of simple, atomic data types combined using constructs such as products (tuples or records) and sums (variants or unions). This compositional nature allows for the creation of complex, hierarchical data structures that can accurately capture the intricate details of molecular structures. By defining molecules using ADTs, we inherently specify the allowed components and their valid combinations, effectively creating a grammar for the representation. This grammar ensures that only semantically meaningful structures can be constructed, preventing the formation of invalid or nonsensical molecules. Furthermore algebraic data types can be made instances of groups, allowing for a uniform, verifiable language to describe rotational symmetries of molecules.
One perk of this representation is the ability to be more general than a Lewis structure as it allows “individual electrons” to be in multiple covalent bonds. This is achieved with a representation published in 1995 [Dietz, 1995]. This allows for modelling delocalised electrons thus allowing for resonance structure, organometallics and multi-center or electron deficient bonds. This is advocated for in recent papers such as those discussing the future of molecular representations[Krenn et al., 2022]. Modelling organometallics has already proved useful in the drug design process, with Platinum based drugs being used to prevent cancer [dilruba2016platinum]. One result of this is the ability to calculate fractional and zero bond orders, often impossible in other common representations. Furthermore, the code creates a data structure to model shells, subshells, and orbitals. The Shell data type represents a shell with a principal quantum number and optional subshells ($s$, $p$, $d$, $f$). Each subshell is represented as a list of orbitals. The Orbital data type represents an individual orbitals with its type, electron count, and orientation. The more general Shells type is defined as a list of Shell, representing the complete electronic structure of an atom.
The ADT introduced is fairly extensible: removing the need for completely different syntaxes to be created for each different kind of molecule. Haskell’s record syntax further enables access of only parts of the molecule needed for a task e.g. in machine learning. This means redundant data is discarded, making machine learning models less wasteful, thus more efficient and interpretable. This is difficult to do with some representations, where dissecting or parsing the string is difficult, so the whole molecule is fed in, including redundant data.
The representation is easily extensbile allowing chemists to adjust their data structure straightforwardly in line with their needs.
[…]
Probabilistic programming languages (PPLs) have emerged as a powerful paradigm for performing Bayesian inference, enabling the representation and manipulation of probability distributions over complex data structures. PPLs provide a unified framework for specifying probabilistic models, performing inference, and making predictions based on observed data. The program calculates the posterior provided a model and an inference strategy, where inference is broadly defined by Gordon et al. [Gordon et al., 2014] as the problem of computing an explicit probability distribution implicit in the program. The idea that a program can represent a probabilistic model goes back to 1978 [Saheb-Djahromi, 1978] and programming languages for Bayesian modelling have existed since at least 1994 [Gilks et al., 1994]. The goal of probabilistic programming as mentioned by [Gordon et al., 2014] is to hide inference inside the compiler to allow the programmer to focus on writing models and have inference automatically done for them.
@@ 2.6 Introduction to the proposal
Further efficiency guarantees are provided by using a lazy language. Haskell as a lazy language, means that it only evaluates functions when needed. This can provide efficiency improvements when simulations are performed. This is very helpful when sampling from molecular orbitals which are represented naturally in this library by higher order functions over random samples. In particular, combining it with LazyPPL [Paquet and Staton, 2021], a recently developed lazy probabilistic programming language that only samples when necessary may provide efficiency improvements.
3 Algebraic Data Types
An algebraic data type (ADT) in Haskell represents data with additional structure and they differentiate functions from data. An ADT provides a structured way to organize data by utilizing type constructors, such as sums and products, to create complex types from simpler ones. This structured organization of data adheres to algebraic laws. Typeclasses in Haskell exemplify this capability by denoting which data types support specific operations and specification of the equational laws that these operations must satisfy can be stated. By making a data type an instance of a typeclass, we declare that the data type supports the operations defined by the typeclass and claim that these operations adhere to the specified equational laws. This allows for the representation of complex algebraic structures, such as algebraic groups within the type system. Moreover, typeclasses enable a high level of abstraction and code reuse, promoting modularity and extensibility in the context of molecular representation, where they can be used to define the algebraic structures and operations relevant to molecules, such as groups for rotational symmetries.

### v2

@@ 3.15 Molecules and programming languages
Inductive Logic Programming (ILP), has seen usage in representing basic molecules [37], but the focus was doing rule induction over graphical data structures, with molecules being an example, rather than presenting ILP as a representational format for cheminformaticians. It is perhaps for this reason that it has had little uptake within the cheminformatics community.
Molecular representations using Inductive Logic Programming (ILP) have been explored in previous research, such as the work by Srinivasan et al. [37]. However, the application of machine learning techniques in conjunction with ILP-based molecular representations has not been extensively discussed in this early work. This lack of emphasis on machine learning aspects may explain the absence of the paper from recent reviews of molecular representations, such as the comprehensive survey by Wigh et al. [12].
In 2018, de Meent et al. [38] argue that the recent successes of artificial intelligence have been caused by the emergence of programming language tools, such as NumPy, that are able to automatically engage in the “tedious” derivation and calculation of gradients. Probabilistic programming languages build on this idea to provide a generic way to write arbitrary models and then calculate the uncertainty of parameter choices.
In the approach presented by Srinivasan et al. [37], molecules are represented as logical rules using an extension of the Prolog programming language. Prolog is a declarative language that enables the expression of logical relationships between variables using first-order logic. Programs written in Prolog consist of a complex network of facts and rules, which are constructed using logical primitives such as conjunction, disjunction, and universal quantification. However, the data types supported by Prolog are limited to lists, integers, floats, and strings, with minimal support for user-defined data types.
@@
The specific extension of Prolog used in [37] is called Progol, which employs inductive logic programming techniques to derive rules that describe the given data. It is worth clarifying that Progol should not be confused with ProbLog, a probabilistic logic programming language that focuses on modeling the probabilities of examples being true based on the probabilities of the underlying relations.
@@
While the work by Srinivasan et al. [37] demonstrates the feasibility of representing molecules using logical rules in an ILP framework, the integration of machine learning techniques with these representations was not explored.
This makes the proposal here the first to suggest to cheminformaticians to use constructs rooted in functional programming languages to be directly involved as representation. The notion that “molecular programming languages” could be possible, occurred in an article about the future of SELFIES [1]. However, the sense attached to that project has been questioned previous to this.
@@
Investigations of type safe functional programming for cheminformatics has occurred before. Recent work has suggested the use of OCaml in writing software for cheminformatics [39, 40]. However, the emphasis in both of these articles as well as the Haskell library Ouch [41], was not on molecular representations but more a tutorial on functional programming. This is highlighted by the papers introduction of lengthy tutorials on functional programming. Furthermore they do not show how Bayesian Inference can be performed on drug structures over algebraic data types which is essential in motivating efficiency improvements in not requiring exploration of syntactically nonsensical regions of molecular space.
4 Probabilistic programming
As discussed in a recent thesis on modular probabilistic programming, and I quote from it through this and the next paragraph [42], the Bayesian approach to probabilistic reasoning is known as Bayesian Inference, and is an increasingly popular method to perform machine learning that enjoys strong mathematical foundations. Bayesian Inference has useful properties, compared to deep learning based machine learning, such as interpretability, explainability and expressiveness with respect to various forms of uncertainty [43, 44]. At the heart of the Bayesian paradigm is the notion of a generative model. Generative models can account for the expressiveness of uncertainty and interpretability. In order to perform probabilistic modelling, knowledge about a domain and the uncertainty associated with that knowledge is encoded into a generative model, using expert knowledge. Inference algorithms estimate parameters of the model, that generate the observed data, by conditioning models on associated observed data. For instance, a climate researcher may create a basic model of a city when studying urban heat islands. The climate researcher would condition the model of the city with empirical data, such as the observed temperatures of the city, in order to infer values of interest such as latent city temperatures or factors that influence the temperature of a city. Probabilistic programming languages provide a generic framework to perform inference.
To highlight the importance of probabilistic programming languages, large multinational technology companies such as Microsoft, Google and Uber have recently started to invest in their own probabilistic programming languages (Infer.NET [45], Edward [46] and Pyro [47] respectively). Increasingly, machine learning use cases find their way into areas as diverse as critical defense national infrastructure [48][49] or medical diagnostics [50]. Society is therefore increasingly in a position that relies on programmers to produce correct machine learning related code. As probabilistic programming becomes more important, the code that underlies probabilistic programming languages will face increasing scrutiny. Fortunately, there are two important programming language research goals which aim to ensure the correctness of code; modularity and type-safety. In particular, they ensure that programmers can be certain early on – at compile time – that composing code will not cause programs to crash or introduce unintended (inference) algorithmic bias. These research goals have two effects; as programming libraries become safer and more modular, machine learning models will become cleaner, simpler and easier to maintain, which will additionally make machine learning more trusted, and perhaps more widely adopted.
@@ 4 Probabilistic programming
5 Methodology: Laziness and The Algebraic Data Type.
An Algebraic Data Type (ADT) is a composite data structure formed by combining simpler data types using sum types (representing alternatives) and product types (representing combinations), which obey algebraic laws [51]. It is a fundamental concept in functional programming, with links to universal algebra, that provide a way to define complex, hierarchical, or (iso)recursive structures, first introduced in 1980 with HOPE [52, 51]. They define the data structure, and its type with which typeclasses then define the operations that can be meaningfully performed on that data, through instantiation.
Our molecular Algebraic Data Type is written in Haskell[53], a lazy, purely functional, type-safe language, with type classes and algebraic data types.
In cheminformatics, we often deal with large datasets of chemical structures (e.g. millions of molecules) and sometimes expensive analyses (e.g. molecular property calculations, conformer generation, docking runs). A lazy programming language like Haskell can help manage these challenges by delaying computations until they are truly needed, and by automatically reusing results once they have been computed.
In many programming languages, as soon as you write an expression—for example, a function call or a chemical property calculation—the language runtime will typically evaluate it immediately. In a lazy language, however, expressions are turned into thunks: little bundles of code waiting to be executed. The language defers (delays) performing the actual computation until it is absolutely required, such as when the result is needed to proceed with another step in the program.
Suppose we define a huge list of candidate molecules, each annotated with an expensive-to-calculate property (e.g. predicted solubility). In a strict language, as soon as we construct this list, the program might try to calculate the property for every molecule upfront, even if we only want the first few results. In a lazy language, by contrast, the expensive property calculations are wrapped up in thunks. These thunks remain unevaluated until you actually iterate over the list and access the relevant element.
Thus, if you only access the first 10 molecules in the dataset, only 10 calculations take place; the remaining 999,990 molecular properties remain as unevaluated thunks.
A second important aspect of laziness is sharing:
once a thunk has been computed, its result is cached (or memoized) so that future accesses do not repeat the same work. If multiple parts of your cheminformatics workflow depend on the same expensive molecular descriptor, the lazy evaluation mechanism ensures that once you compute it once, the result is re-used everywhere else automatically.
  • On-Demand Computation: We might have large pipelines combining multiple steps (e.g. filtering by molecular weight, docking simulations, then clustering). If we never end up needing a particular property, we waste no effort computing it in a lazy setting.
  • Improved Compositionality: With lazy evaluation, we can write “infinite” or “very large” data streams and process them in a streaming fashion without ever fully materializing them in memory. This is particularly helpful when investigating structure-generation procedures or enumerations in a conceptually infinite chemical space, since only the portion of the stream you consume is actually computed.
  • Parallelization and Caching: Lazy languages can simplify certain parallelization strategies since once a thunk is forced in one thread, its result can be shared across other threads (in purely functional languages like Haskell).
Our Algebraic Data Type represents molecular constitution with the Dietz representation, three dimensional chemical information via atomic coordinate data, and provides a novel representation of quantum information via the Orbital type. Through the use of Haskell’s type classes, we specify allowed components and combinations of molecules, ensuring only semantically meaningful structures are represented. Additionally, algebraic data types can be shown to be a group (or a monoid etc), meaning rotational symmetry of molecules can be described.
By using the Dietz representation’s general model for bonding systems, the ADT is inherently able to represent delocalised bonding in general, without extensive labelling of bond types. Resonance structures, organometallics, multi-center and electron deficient bonds may all be represented. Fractional, and zero bond orders may also be calculated.
Using atomic coordinate data, all molecules with stereoisomerism may be represented and distinguished, and dihedral bond angles, torsion angles, and bond lengths may be calculated. Such information adds minimally to the space complexity, but yet can be crucial when predicting physical propreties of drugs.
The Dietz represenation of constitution, combined with atomic coordinate data then solves the three identified limitations of SELFIES for going beyond organic chemistry[1]: the ability to represent delocalised bonding, explicit hydrogen representation, and the ability to represent complex stereochemistry.
The novel orbital data type creates a data structure to model shells, subshells, and orbitals, completely representing the electronic structure of an atom. Quantum information can then be used for machine learning applications, and for quantum chemical calculations in Molecular Orbital theory, as can be required with organometallics.
Apart from the obvious benefits of being able to represent molecules which cannot currently be represented using string-based representations (even with the creation of new and complicated grammars), our ADT confers a crucial benefit. By directly synthesizing molecules in accordance with the grammar defined by the ADT, machine learning processes can operate directly on the program or grammar itself, without the need for external context or to encode the representation into another format.
5.1 Haskell
Haskell’s type classes denote which data types support specific operations, and in some cases can also specify laws that the operations should satisfy. By making data types an instance of a type class, we specify that the data type supports the defined operations, and ensure the operations adhere to the operational laws.
What is novel is the integration of probabilistic programming with Haskell into a concrete use case. Probabilistic programming languages themselves in haskell have been created before[54, 55, 42], laying the groundwork for usage with chemical structures. The existence of libraries such as LazyPPL[56] (demonstrated later in the paper), makes the ADT possible to use with Bayesian machine learning applications, such as drug design, which have been attempted before [57, 58, 54], and is an emerging area of research. The “lazy” (non strict, or call-by-need) property of Haskell, as discussed earlier, can provide efficiency, for instance when sampling from expensive functions or performing simulations only when necessary.
The record syntax in Haskell also enables access of only parts of the molecule needed for a task , e.g. in machine learning. Non-necessary data is not accessed, making machine learning models potentially less wasteful. This is difficult to do with non-local encoding of information, such as with SMILES or SELFIES strings, where a whole molecule, as a string, must be parsed.